# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v4

@@ 4. Inference on Estimands
After fixing the estimator—and hence a statistical functional $\theta^{*}(P)$ defined on $\mathcal{P}^{*}$—we naturally seek to quantify sampling uncertainty, but misspecification can complicate inference. Under misspecification, standard error formulas derived under $\mathcal{P}$ may no longer be correct; corresponding confidence intervals may undercover the functional $\theta^{*}(P)$ when $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. This is a well-known motivation for misspecification-robust standard errors for maximum likelihood (White, 1982).
Similar results exist for GMM (Imbens, 1997; Hall and Inoue, 2003; Hansen and Lee, 2021), yet they are seldom used in practice (analytic misspecification-robust standard errors do not appear to be used in any of the papers in our GMM sample). Since these results seem to be less well-known, we sketch an asymptotic expansion illustrating how misspecification impacts GMM, as well as the intuition for robust standard errors. Let $\hat{\psi}_{\Omega}$ be the GMM estimator with weighting matrix $\hat{\Omega}=\Omega+o_{P}(1)$ and $\psi^{*}_{\Omega}$ be its corresponding statistical functional. [^19] Since we consider behavior at a fixed $P,$ where possible we suppress the dependence on $P$ to simplify notation. Under statistically correct specification, $\psi_{\Omega}^{*}$ is simply the GMM parameter defined by $\mathbb{E}_{P}[g(X_{i},\psi_{\Omega}^{*})]=0$ (assumed unique). Under misspecification, $\psi_{\Omega}^{*}$ minimizes the population GMM criterion $\phi\mapsto(\mathbb{E}_{P}[g(X_{i},\phi)])^{\prime}\Omega\mathbb{E}_{P}[g(X_{i},\phi)]$. Let $\Gamma=\mathbb{E}_{P}[\nabla g(X_{i},\psi_{\Omega}^{*})]$ and
@@
By the first-order conditions for GMM, $\hat{\Gamma}^{\prime}\hat{\Omega}g_{n}(\hat{\psi}_{\Omega})=0=\Gamma^{\prime}\Omega\mathbb{E}_{P}[g_{n}(\psi_{\Omega}^{*})]$.
To derive the standard errors for the GMM estimator under misspecification, note that by a first-order Taylor expansion and $\sqrt{n}$ consistency of $\hat{\psi}_{\Omega},$
@@
$=\sqrt{n}g_{n}(\psi_{\Omega}^{*})+\Gamma\sqrt{n}(\hat{\psi}_{\Omega}-\psi_{\Omega}^{*})+o_{P}(1).$ (2)
@@
$(\hat{\Gamma}^{\prime}\hat{\Omega}\Gamma)\sqrt{n}(\hat{\psi}_{\Omega}-\psi_{\Omega}^{*})=\sqrt{n}\Gamma^{\prime}\Omega g_{n}(\psi_{\Omega}^{*})+\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})+o_{P}(1).$ (3)
@@
where $\Sigma(P)\equiv\text{aVar}_{P}(\sqrt{n}g_{n}(\psi_{\Omega}^{*}))$, and $\text{aVar}_{P}(\cdot)$ denotes the asymptotic variance under $P$. However, when the model is statistically misspecified $\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})$ is non-negligible and, under regularity conditions, converges to a Gaussian with nonzero, consistently estimable variance. Thus, the familiar sandwich formula is no longer correct and misspecification-robust standard errors should be used instead.
@@
Since many economic models are statistically misspecified, we are skeptical that the conventional GMM formula accurately captures the sampling variability of GMM estimators in practice. Parallel to White (1982)’s widely-adopted suggestion for maximum likelihood, we recommend that robust standard errors be used for GMM as well. Again parallel to maximum likelihood, robust standard errors for GMM do no harm, at least in large samples: They are equal to the conventional standard errors when $P$ happens to be well-specified, but remain valid for inference on $\psi_{\Omega}^{*}(P)$ when $P\not\in\mathcal{P}$. Consequently, provided $\vartheta(\psi)$ is smooth in $\psi,$ delta-method standard errors will be valid for inference on $\theta^{*}_{\Omega}(P)=\vartheta(\psi_{\Omega}^{*}(P))$.
@@
The robust standard error formula (e.g. Hall and Inoue, 2003) depends on additional terms not present in the usual formula, and can be complicated in some cases though software implementations are available. [^20] For instance in the dcivreg command in Stata, which implements the misspecification-robust approach of Hwang et al. (2022) for linear IV. An alternative for settings with iid data is the standard nonparametric bootstrap, which would approximate the limit distribution in (3) when $\theta^{*}(P)$ varies smoothly with PP. [^21] As discussed by Hong and Li (2025), the smoothness requirements for $\sqrt{n}$ asymptotic normality of GMM under misspecification are stronger than those for the correctly specified case. The authors provide a bootstrap which is robust to insufficient smoothness. For bootstrapping GMM models, some propose recentering the bootstrapped moments so as to enforce the moment restriction in the “bootstrap world” (Brown and Newey 2002; also see Section 3.7 in Horowitz 2001). When combined with a t-ratio bootstrap, this approach generates asymptotic refinements under statistically correct specification. Unfortunately, however, recentering yields invalid inference under statistical misspecification. Intuitively, recentering imposes correct specification in the bootstrap data generating process. Thus, the asymptotic variance of the estimator under the recentered bootstrap matches the conventional (invalid) sandwich formula. [^22] Similarly, bootstrapping the moments but using the conventional formula to approximate the estimator variance, as done in some papers in our GMM sample, yields non-robust standard errors. We thus recommend against recentering moments.
@@
While the asymptotic expansion (3) is valid under misspecification it, and the resulting standard errors, still relies on regularity conditions, such as the existence of a unique interior solution $\psi^{*}_{\Omega}(P)$ in the population GMM problem, strong identification of $\psi^{*}_{\Omega}(P)$, and differentiability of $\vartheta$. Even when these conditions fail, however, the nonparametric bootstrap distribution for the estimate $\hat{\theta}_{\Omega}$ is asymptotically equivalent to the Bayesian posterior distribution for the estimand $\theta_{\Omega}^{*}(P)$ under a particular nonparametric and non-informative limit of priors. [^23] See e.g. Andrews and Shapiro (2024) for discussion of the relevant priors, and Appendix D of that paper for sufficient conditions for asymptotic equivalence. If we use the Bayes rather than nonparametric bootstrap, then the equivalence is exact. Thus, if we use bootstrap percentile intervals for inference on $\theta_{\Omega}^{*},$ they offer correct large-sample coverage for well-behaved estimands $\theta^{*}_{\Omega}(P)$ regardless of model misspecification, while also providing a “fallback” interpretation as a Bayesian credible set for $\theta^{*}_{\Omega}(P)$ in less well-behaved cases. The bootstrap also provides a natural tool for diagnosing the failure of the conventional, non-misspecification-robust GMM standard errors: if the bootstrap distribution appears approximately normal and centered at $\hat{\theta}_{\Omega},$ but differs from the normal suggested by the conventional standard error, this suggests the expansion (3) may be reasonable but that the conventional, non-misspecification-robust asymptotic approximation is not. See Andrews and Shapiro (2024) for formal results on this approach to evaluating asymptotic approximations. $\blacksquare$
@@
While White (1980) is primarily framed around heteroskedasticity robustness, these standard errors for linear regression are also robust to nonlinearity of the conditional mean function and so represent a very widely used application of the misspecification-robust approach. More recently, Lee (2018) highlights the invalidity of conventional GMM standard errors in two stage least squares with heterogeneous effects and multiple instruments, and specializes the formula of Hall and Inoue (2003) to this case. Lee (2014) shows that bootstrapping a misspecification-robust t-statistic yields asymptotic refinements for inference on the GMM estimand, offering an alternative to recentering, and Hansen and Lee (2021) derive misspecification-robust standard errors for the iterated-to-convergence GMM estimator. Beyond standard errors, misspecification also invalidates many identification-robust procedures, and Kleibergen and Zhan (2025) derive an identification- and misspecification-robust confidence interval for the continuously updating GMM estimand. Finally, Müller (2013) considers Bayesian decision making in misspecified models and shows that for decision problems concerning ML estimands, decision rules based on an artificial posterior constructed using robust standard errors asymptotically dominate conventional Bayesian inference. $\blacksquare$
@@ 5. Local Misspecification
Our study of statistical properties (e.g. probability limits and standard errors) has so far focused on fixed, or “global” misspecification, since we have held $P$ fixed and considered behavior as the sample size grows. As $n$ grows, however, any fixed distribution $P\notin\mathcal{P}$ becomes detectable with probability one. In a GMM setting, for instance, J-statistics
@@
diverge to infinity under global misspecification, rather than to their limiting $\chi^{2}_{k-p}$ distribution under statistically correct specification (Hansen, 1982).
In many economic applications the degree of statistical misspecification appears to be roughly on the same order as statistical uncertainty. For instance, J-statistics, while larger than expected under correct specification, are not always vastly so. Estimates from different weighting matrices may be statistically distinguishable from one another, yet numerically and substantively fairly close. Motivated by this observation, and following Newey (1985), a growing econometrics literature studies models of local misspecification, which assume that the sample of size $n$ is generated by a data-generating process $P_{n}$ such that at some fixed value $\psi$ for the GMM parameters,
$\mathbb{E}_{P_{n}}\left[{g(X_{i},{\psi})}\right]=\frac{\eta}{\sqrt{n}}.$
@@
As with other local asymptotic approximations, for instance in the literature studying weak instruments (Staiger and Stock, 1997), we do not imagine that the data distribution would mechanically change were we able to gather more data. Instead, local misspecification is a modeling device to ensure that our asymptotic approximation reflects a salient feature of many finite-sample problems, namely that the apparent degree of misspecification is of the same order as sampling uncertainty. Smaller misspecification would vanish compared to sampling noise; larger misspecification would swamp—and invalidate, as discussed in the last section—standard errors. Hence local misspecification is the appropriate regime for studying trade-offs between estimand choice and estimator precision.
Local misspecification also allows important analytical simplifications. Specifically, in locally misspecified settings many statistics behave like their correctly-specified analogs, save for the addition of extra bias terms. In particular, Newey (1985) Theorem 4 shows that under local misspecification the GMM estimate is asymptotically linear in the moments evaluated at $\psi,$
$\hat{\psi}_{\Omega}-\psi=-(\Gamma^{\prime}\Omega\Gamma)^{-1}\Gamma^{\prime}\Omega g_{n}(\psi)+o_{p}\left(\frac{1}{\sqrt{n}}\right)$ (4)
@@
The local misspecification framework has been used in a number of recent papers. For instance, Andrews et al. (2017) suggest researchers looking to give intuition for the behavior of their estimates may report estimates of $(\Gamma^{\prime}\Omega\Gamma)^{-1}\Gamma^{\prime}\Omega$, since this describes how misspecification in the moments (i.e. different values of $\eta$) feeds forward into bias in the estimates. Armstrong and Kolesár (2021) impose the further restriction that $\theta=\vartheta(\psi),$ so the GMM model is econometrically correctly specified in the case where $\eta=0,$ and derive optimal confidence intervals for $\theta$ under restrictions on the possible values for $\eta.$ Bonhomme and Weidner (2022) consider models with additional structure where some model components may be locally misspecified, and derive asymptotically optimal estimators under bounds on the degree of misspecification. Similarly, for a class of moment-condition models imposing parametric assumptions Christensen and Connault (2023) derive bounds on quantities of interest under both local and global misspecification.
Despite this recent activity using the local misspecification framework, there remains scope for new and interesting results in this area. To illustrate, in the remainder of this section we use the local misspecification framework to develop what is to the best of our knowledge a new result, characterizing the relationship between the J-statistic and the class of GMM estimates obtained by varying the weighting matrix.
@@ 5.1. J-Statistics, not J-Tests
Relatively few economics papers report J-statistics. In our GMM sample, for instance, only 3 out of the 22 over-identified papers report J-statistics, either in the paper or the online appendix. [^24] One additional paper, Mueller et al. (2021) reports the value of the two-step SMM objective function at its minimum. Because they use the optimal weighting matrix (estimated by bootstrap), this objective value is equal to the J-statistic absent simulation error. One reason may be that few economists expect their models to hold exactly. Under that view, a non-rejection signals insufficient power of the J-test, not correct specification. This suggests a focus on “economic” rather than statistical measures of model fit. Consistent with this, 19 of the 22 over-identified papers in our GMM sample report some version of an informal identification check, often by plotting the model-implied values of a set of moments against either (i) the data moments that were used to estimate the model or (ii) a “held out” set of moments not used in estimation, which may be interpreted as estimation moments assigned zero weight by $\hat{\Omega}$. Both of these approaches convey information about the extent to which the model’s over-identifying restrictions appear to hold in the sample, but lack the formal statistical interpretation of the J-statistic.
@@
To better understand the behavior of J statistics in practice, we attempted to calculate J statistics for the papers in our GMM sample which did not already report them. In this exercise we were limited to those papers with accessible data, and where the optimally-weighted GMM objective was sufficiently well-behaved that we were confident in the convergence of our numerical optimization routines. Subject to these restrictions, we were able to compute J-statistics for an additional three papers beyond those which initially reported them. Table 1 collects the J-statistics for these papers, along with the original three. Strikingly, for the 6 papers in our sample where we were able to obtain J statistics, in only one case is the model rejected at the 5% level (although three papers have specifications rejected at the 10% level). Also striking, while the number of papers is limited, there is not obvious selection in which papers report J-statistic. Indeed, the papers for which we computed J-statistics have, on the whole, larger p-values (and thus, in that sense, less evidence of misspecification) than the papers which originally reported J-statistics.
Table 1. $X$
Paper Model J-statistic dF p-value Reported in the paper?
Table 4, Column 1 3.05 1 0.08
Abebe et al. (2021b) Table 4, Column 2 4.67 3 0.20 No
Table 4, Column 4 4.28 2 0.12
Bourreau et al. (2021) Table 4, Column 2 50.88 7 0.00 J-statistic and dF
Mueller et al. (2021) Table 5 0.462 3 0.93 No
García-Jimeno et al. (2022) Table 2, Column 2 2.16 5 0.83 J-statistic and p-value
Table 6, Column 7 0.361 1 0.55
Karahan et al. (2024) Table 7, Column 4 5.813 2 0.05 p-value
Table 7, Column 5 0.327 1 0.57
Kelley et al. (2024) Table 7 0.005 1 0.94 No
  • Notes: This table reports J-statistics and associated degrees of freedom and p-values for six papers in our “GMM sample.” See the text for inclusion criteria.
To develop our result, we consider classes of weighting matrices with minimum and maximum eigenvalue bounded below and above by $\kappa^{-1}$ and $\kappa,$ respectively, for $\kappa\in[1,\infty)$ a positive constant:
@@
Provided the linear approximation (4) holds uniformly over such classes as $n\to\infty,$ the J-statistic characterizes, in a precise sense, the extent to which GMM results vary across weighting matrix choices. [^25] We limit attention to $\Omega\in\mathcal{W}_{\kappa}$ because there is no hope that the linear approximation is uniformly valid over classes that whose closure includes e.g. weight matrices with severely reduced rank.
@@
as $n\to\infty$ for all fixed $\kappa>0.$ Moreover, we have estimators $\hat{\Gamma}_{\Omega}$ and $\hat{\Sigma}_{\Omega}$ such that $\hat{\Gamma}_{\Omega}\overset{p}{\longrightarrow}\Gamma$ and $\hat{\Sigma}_{\Omega}\overset{p}{\longrightarrow}\Sigma$ uniformly over $\mathcal{W}_{\kappa}$, where both limits are full rank and do not depend on $\Omega,$ and $\sqrt{n}\lVert g_{n}(\hat{\psi}_{\Sigma^{-1}})-g_{n}(\psi)+\Gamma(\hat{\psi}_{\Sigma^{-1}}-\psi)\rVert\overset{p}{\longrightarrow}0.$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left(\sqrt{n}\cdot d_{H}\left(\hat{\Theta}\left(\kappa,\tau\right),\hat{\Theta}(\tau)\right)>\varepsilon\right)=0.$
@@
Under the assumptions of 5.2, consider the t-statistic for testing $H_{0}:\theta=\theta_{0}$ based on the GMM estimator with weighting matrix $\Omega$:

### v5

@@ 4. Inference on Estimands
After fixing the estimator—and hence a statistical functional $\theta^{*}(P)$ defined on $\mathcal{P}^{*}$—we naturally seek to quantify sampling uncertainty, but misspecification can complicate inference. Under misspecification, standard error formulas derived under $\mathcal{P}$ may no longer be correct; corresponding confidence intervals may undercover the functional $\theta^{*}(P)$ when $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. This is a well-known motivation for misspecification-robust standard errors for maximum likelihood (white1982maximum).
Similar results exist for GMM (imbens1997one; hall2003large; hansen2021inference), yet they are seldom used in practice (analytic misspecification-robust standard errors do not appear to be used in any of the papers in our GMM sample). Since these results seem to be less well-known, we sketch an asymptotic expansion illustrating how misspecification impacts GMM, as well as the intuition for robust standard errors. Let $\hat{\psi}_{\Omega}$ be the GMM estimator with weighting matrix $\hat{\Omega}=\Omega+o_{P}(1)$ and $\psi^{*}_{\Omega}$ be its corresponding statistical functional. [^21] Since we consider behavior at a fixed $P,$ where possible we suppress the dependence on $P$ to simplify notation. Under statistically correct specification, $\psi_{\Omega}^{*}$ is simply the GMM parameter defined by $E_{P}[g(X_{i},\psi_{\Omega}^{*})]=0$ (assumed unique). Under misspecification, $\psi_{\Omega}^{*}$ minimizes the population GMM criterion $\phi\mapsto(E_{P}[g(X_{i},\phi)])^{\prime}\Omega E_{P}[g(X_{i},\phi)]$. Let $\Gamma=E_{P}[\nabla g(X_{i},\psi_{\Omega}^{*})]$ and
@@
By the first-order conditions for GMM, $\hat{\Gamma}^{\prime}\hat{\Omega}g_{n}(\hat{\psi}_{\Omega})=0=\Gamma^{\prime}\Omega E_{P}[g_{n}(\psi_{\Omega}^{*})]$.
To derive the standard errors for the GMM estimator under misspecification, note that by a first-order Taylor expansion and $\sqrt{n}$-consistency of $\hat{\psi}_{\Omega},$
@@
(2) $=\sqrt{n}g_{n}(\psi_{\Omega}^{*})+\Gamma\sqrt{n}(\hat{\psi}_{\Omega}-\psi_{\Omega}^{*})+o_{P}(1).$
@@
(3) $(\hat{\Gamma}^{\prime}\hat{\Omega}\Gamma)\sqrt{n}(\hat{\psi}_{\Omega}-\psi_{\Omega}^{*})=\sqrt{n}\Gamma^{\prime}\Omega g_{n}(\psi_{\Omega}^{*})+\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})+o_{P}(1).$
@@
where $\Sigma(P)\equiv\text{aVar}_{P}(\sqrt{n}g_{n}(\psi_{\Omega}^{*}))$ for $\text{aVar}_{P}(\cdot)$ the asymptotic variance under $P$. However, when the model is statistically misspecified $\sqrt{n}(\hat{\Gamma}^{\prime}\hat{\Omega}-\Gamma^{\prime}\Omega)g_{n}(\psi_{\Omega}^{*})$ is non-negligible and, under regularity conditions, converges to a mean-zero Gaussian with nonzero, consistently estimable variance. Thus, the familiar sandwich formula is no longer correct and misspecification-robust standard errors should be used instead.
@@
Since many economic models are statistically misspecified, it is questionable whether the conventional GMM formula accurately captures the sampling variability of GMM estimators in practice. Parallel to white1982maximum’s widely-adopted suggestion for maximum likelihood, we recommend that robust standard errors be used for GMM as well. Again parallel to maximum likelihood, robust standard errors for GMM do no harm, at least in large samples: They are equal to the conventional standard errors when the model happens to be statistically well-specified, but remain valid for inference on $\psi_{\Omega}^{*}(P)$ when $P\not\in\mathcal{P}$. Consequently, provided $\vartheta(\psi)$ is smooth in $\psi,$ delta-method standard errors will be valid for inference on $\theta^{*}_{\Omega}(P)=\vartheta(\psi_{\Omega}^{*}(P))$.
@@
The robust standard error formula (e.g. hall2003large) depends on additional terms not present in the usual formula, and can be complicated in some cases though software implementations are available. [^22] For instance in the dcivreg command in Stata, which implements the misspecification-robust approach of Hwangetal2022 for linear IV. An alternative for settings with i.i.d. data is the standard nonparametric bootstrap, which approximates the limit distribution in (3) when $\theta^{*}(P)$ varies smoothly with PP. [^23] As discussed by Hong_Li_2025, the smoothness requirements for $\sqrt{n}$ asymptotic normality of GMM under misspecification are stronger than those for the correctly specified case. The authors provide a bootstrap which is robust to insufficient smoothness. For bootstrapping GMM models, some propose recentering the bootstrapped moments so as to enforce the moment restriction in the “bootstrap world” (brown2002generalized; also see Section 3.7 in horowitz2001bootstrap). When combined with a t-ratio bootstrap, this approach generates asymptotic refinements under statistically correct specification. Unfortunately, however, recentering yields invalid inference under statistical misspecification. Intuitively, recentering imposes correct specification in the bootstrap data generating process. Thus, the asymptotic variance of the estimator under the recentered bootstrap matches the conventional (invalid) sandwich formula. [^24] Similarly, bootstrapping the moments but using the conventional formula to approximate the estimator variance, as done in some papers in our GMM sample, yields non-robust standard errors. We thus recommend against recentering moments.
@@
While the asymptotic expansion (3) is valid under misspecification, both the expansion itself and the resulting standard errors rely on regularity conditions, including the existence of a unique interior solution $\psi^{*}_{\Omega}(P)$ in the population GMM problem, strong identification of $\psi^{*}_{\Omega}(P)$, and differentiability of $\vartheta$. Even when these conditions fail, however, the nonparametric bootstrap distribution for the estimate $\hat{\theta}_{\Omega}$ is asymptotically equivalent to the Bayesian posterior distribution for the estimand $\theta_{\Omega}^{*}(P)$ under a particular nonparametric and non-informative limit of priors. [^25] See e.g. Chamberlain01012003; andrewsshapiro2024bootstrap for discussion of the relevant priors, and Appendix D of the latter paper for sufficient conditions for asymptotic equivalence. If we use the Bayes rather than nonparametric bootstrap, then the equivalence is exact. Thus, if we use bootstrap percentile intervals for inference on $\theta_{\Omega}^{*},$ they offer correct large-sample coverage for well-behaved estimands $\theta^{*}_{\Omega}(P)$ regardless of model misspecification, while also providing a “fallback” interpretation as a Bayesian credible set for $\theta^{*}_{\Omega}(P)$ in less well-behaved cases. The bootstrap also provides a natural tool for diagnosing the failure of the conventional, non-misspecification-robust GMM standard errors: if the bootstrap distribution appears approximately normal and centered at $\hat{\theta}_{\Omega},$ but differs from the normal suggested by the conventional standard error, this suggests the expansion (3) may be reasonable but that the conventional, non-misspecification-robust asymptotic approximation is not. See andrewsshapiro2024bootstrap for formal results on this approach to evaluating asymptotic approximations. $\blacksquare$
@@
While white1980robust is primarily framed around heteroskedasticity robustness, these standard errors for linear regression are also robust to nonlinearity of the conditional mean function and so represent a very widely used application of the misspecification-robust approach. More recently, Lee2018tsls highlights the invalidity of conventional GMM standard errors in two stage least squares with heterogeneous effects and multiple instruments, and specializes the formula of hall2003large to this case. lee2014bootstrap shows that bootstrapping a misspecification-robust t-statistic yields asymptotic refinements for inference on the GMM estimand, offering an alternative to recentering, and hansen2021inference derive misspecification-robust standard errors for the iterated-to-convergence GMM estimator. Beyond standard errors, misspecification also invalidates many identification-robust procedures: kleibergen2025CUGMM derive an identification- and misspecification-robust confidence interval for the continuously updating GMM estimand. Finally, mueller2013sandwich considers Bayesian decision making in misspecified models and shows that for decision problems concerning maximum likelihood estimands, decision rules based on an artificial posterior constructed using robust standard errors asymptotically dominate conventional Bayesian inference. $\blacksquare$
@@ 5. Local Misspecification
Our study of statistical properties (e.g. probability limits and standard errors) has so far focused on fixed, or “global” misspecification, since we have held $P$ fixed and considered behavior as the sample size grows. As $n$ grows, however, any fixed distribution $P\notin\mathcal{P}$ becomes detectable with probability one under mild conditions. In a GMM setting, for instance, J-statistics
@@
diverge to infinity under global misspecification, rather than to their limiting $\chi^{2}_{k-p}$ distribution under statistically correct specification (hansen1982gmm).
In many economic applications the degree of statistical misspecification appears to be roughly on the same order as statistical uncertainty. For instance, J-statistics, while larger than expected under correct specification, are not always vastly so. Estimates from different weighting matrices may be statistically distinguishable from one another, yet numerically and substantively fairly close. Motivated by this observation, and following newey1985misspec, a growing econometrics literature studies models of local misspecification, which assume that the sample of size $n$ is generated by a data-generating process $P_{n}$ such that at some fixed value $\psi$ for the GMM parameters,
$E_{P_{n}}\left[{g(X_{i},{\psi})}\right]=\frac{\eta}{\sqrt{n}}.$
@@
As with other local asymptotic approximations, for instance in the literature studying weak instruments (staiger1997weakiv), we do not imagine that the data distribution would mechanically change were we able to gather more data. Instead, local misspecification is a modeling device to ensure that our asymptotic approximation reflects a salient feature of many finite-sample problems, namely that the apparent degree of misspecification is of the same order as sampling uncertainty. Smaller misspecification would vanish compared to sampling noise; larger misspecification would swamp—and invalidate, as discussed in the last section—standard errors. Hence local misspecification is the appropriate regime for studying trade-offs between estimand choice and estimator precision.
Local misspecification also allows important analytical simplifications. Specifically, in locally misspecified settings many statistics behave like their correctly-specified analogs, save for the addition of extra bias terms. In particular, (newey1985misspec, Lemma 4) shows that under local misspecification the GMM estimate is asymptotically linear in the moments evaluated at $\psi,$
(4) $\hat{\psi}_{\Omega}-\psi=-(\Gamma^{\prime}\Omega\Gamma)^{-1}\Gamma^{\prime}\Omega g_{n}(\psi)+o_{p}\left(\frac{1}{\sqrt{n}}\right)$
@@
The local misspecification framework has been used in several recent papers. For instance, andrews2017measuring suggest researchers report estimates of $(\Gamma^{\prime}\Omega\Gamma)^{-1}\Gamma^{\prime}\Omega$ to give intuition for the behavior of their estimates, since this describes how misspecification in the moments (i.e. different values of $\eta$) feeds forward into bias in the estimates. armstrong2021sensitivity impose the further restriction that $\theta=\vartheta(\psi),$ so the GMM model is econometrically correctly specified in the case where $\eta=0,$ and derive optimal confidence intervals for $\theta$ under restrictions on the possible values for $\eta.$ bonhomme2022minimizing consider models with additional structure where some model components may be locally misspecified, and derive asymptotically optimal estimators under bounds on the degree of misspecification. Similarly, for a class of moment-condition models imposing parametric assumptions, christensen2023robustness derive bounds on quantities of interest under both local and global misspecification.
Despite this recent activity using the local misspecification framework, there remains scope for new results in this area. To illustrate, in the remainder of this section we use the local misspecification framework to develop—to the best of our knowledge—a new result, characterizing the relationship between the J-statistic and the class of GMM estimates obtained by varying the weighting matrix.
@@ 5.1. J-Statistics, not J-Tests
Relatively few economics papers report J-statistics. In our GMM sample, for instance, only 3 out of the 22 over-identified papers report J-statistics, either in the paper or the online appendix. [^26] One additional paper, mueller2021jobseekers, reports the value of the two-step SMM objective function at its minimum. Because they use the optimal weighting matrix estimated by bootstrap, this objective value is equal to the J-statistic absent simulation error. One reason may be that few economists expect their models to hold exactly. Under that view, a non-rejection signals insufficient power of the J-test, not correct specification. This suggests a focus on “economic” rather than statistical measures of model fit. Consistent with this, 19 of the 22 over-identified papers in our GMM sample report some version of an informal identification check, often by plotting the model-implied values of a set of moments against either (i) the data moments that were used to estimate the model or (ii) a “held out” set of moments not used in estimation, which may be interpreted as estimation moments assigned zero weight by $\hat{\Omega}$. Both of these approaches convey information about the extent to which the model’s over-identifying restrictions appear to hold in the sample, but lack the formal statistical interpretation of the J-statistic.
@@
To better understand the behavior of J-statistics in practice, we attempted to calculate J statistics for the papers in our GMM sample which did not already report them. In this exercise we were limited to those papers with accessible data and code, that reported standard errors, and where we could implement our numerical optimization routines and were confident in their convergence. Subject to these restrictions, we were able to compute J-statistics for an additional three papers beyond those which initially reported them. Table 1 collects the J-statistics for these papers, along with the original three. Strikingly, for the 6 papers in our sample where we were able to obtain J-statistics, in only one case is the model rejected at the 5% level (although three papers have specifications rejected at the 10% level). Also striking, while the number of papers is limited, there is not obvious selection in which papers report J-statistics. Indeed, the papers for which we computed J-statistics have, on the whole, larger p-values (and thus, in that sense, less evidence of misspecification) than the papers which originally reported J-statistics.
Table 1. J-statistics for papers in the “GMM sample”
Paper Model $\sqrt{J}$ dF p-value Reported in the paper?
Table 4, Column 1 1.75 1 0.08
abebe2021selectionoftalent Table 4, Column 2 2.16 3 0.20 No
Table 4, Column 4 2.07 2 0.12
bourreau2021marketentry Table 4, Column 2 7.13 7 0.00 J-statistic and dF
mueller2021jobseekers Table 5 0.68 3 0.93 No
gj2022information Table 2, Column 2 1.47 5 0.83 J-statistic and p-value
Table 6, Column 7 0.60 1 0.55
karahan2024demographic Table 7, Column 4 2.41 2 0.05 p-value
Table 7, Column 5 0.57 1 0.57
kelley2024monitoring Table 7 0.07 1 0.94 No
  • Notes: This table reports the square root of the J-statistics and associated degrees of freedom and p-values for six papers in our “GMM sample.” See the text for inclusion criteria.
To develop our formal result on J-statistics, we consider classes of weighting matrices with minimum and maximum eigenvalue bounded below and above by $\kappa^{-1}$ and $\kappa,$ respectively, for $\kappa\in[1,\infty)$ a positive constant:
@@
Provided the linear approximation (4) holds uniformly over such classes as $n\to\infty,$ the J-statistic characterizes, in a precise sense, the extent to which GMM results vary across weighting matrix choices. [^27] We limit attention to $\Omega\in\mathcal{W}_{\kappa}$ because there is no hope that the linear approximation is uniformly valid over classes whose closure includes e.g. weight matrices with severely reduced rank.
@@
as $n\to\infty$ for all fixed $\kappa\geq 1.$ Moreover, we have estimators $\hat{\Gamma}_{\Omega}$ and $\hat{\Sigma}_{\Omega}$ such that $\hat{\Gamma}_{\Omega}\overset{p}{\longrightarrow}\Gamma$ and $\hat{\Sigma}_{\Omega}\overset{p}{\longrightarrow}\Sigma$ uniformly over $\mathcal{W}_{\kappa}$, where both limits are full rank and do not depend on $\Omega,$ and $\sqrt{n}\lVert g_{n}(\hat{\psi}_{\Sigma^{-1}})-g_{n}(\psi)+\Gamma(\hat{\psi}_{\Sigma^{-1}}-\psi)\rVert\overset{p}{\longrightarrow}0.$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left(\sqrt{n}\cdot d_{H}\left(\hat{\Theta}\left(\kappa,\tau\right),\hat{\Theta}(\tau)\right)>\varepsilon\right)=0.$
@@
Under 5.1, consider the t-statistic for testing $H_{0}:\theta=\theta_{0}$ based on the GMM estimator with weighting matrix $\Omega$: